The Fermi energy level for copper is 6.25 eV. Determine the temperature at 

which there is a 1% probability that an energy state 0.30 eV below the Fermi 

energy level will not contain an electron.


1
Expert's answer
2021-04-01T01:09:23-0400

The probability that the state is empty, is 


"1-f(E_V)=1-\dfrac{1}{1+e^{(E_V-E_F)/(k_BT)}}"

Substitute


"0.01=1-\dfrac{1}{1+e^{(-0.30\ eV)/(k_BT)}}"

"e^{(-0.30\ eV)/(k_BT)}=\dfrac{1}{1-0.01}-1"

"e^{(0.30\ eV)/(k_BT)}=99"

"\dfrac{0.30\ eV}{k_BT}=\ln99"


"T=\dfrac{0.30\ eV}{k_B\ln99}"



"k_B=8.6173\times10^{-5}\ eV\cdot K^{-1}"



"T=\dfrac{0.30\ eV}{8.6173\times10^{-5}\ eV\cdot K^{-1}\ln99}"

"T=757.62\ K"